OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region enclosed by the graphs of y=12-x & y=3x-4, and x-axis about the y-axis.

OpenStudy (tkhunny):

What are your efforts? Is this your first "rotate the region" problem?

OpenStudy (anonymous):

no, we are doing them in calc 2. I have worked the problem out twice already and it says that i am wrong, so i don't know what i am doing wrong or what other ways to do it

OpenStudy (anonymous):

the answer i got was (11392/27)pi

OpenStudy (anonymous):

This looks like applications of integration.

OpenStudy (anonymous):

yeah, i know the application. it's a matter of the process I drew the graphs/region, changed the eqs to x=... because it asks for rotation about the y axis, therefore making it 'dy', and then integrated using the equation: pi S(from 0 to 8) [R^2outer-R^2inner]dy

OpenStudy (tkhunny):

Please show your work. Anything. Rotating around the x-axis is \(\dfrac{3139}{9}\pi\)

OpenStudy (anonymous):

yeah no problem. can i take a phone pic and put on here?

OpenStudy (tkhunny):

Rotating around the x-axis can be either dy or dx. Rotating around the y-axis can be either dx or dy. Don't get lost in one methodology or one way of thinking.

OpenStudy (anonymous):

yes it can, sorry for not clarifying. we are supposed to use disk/shell method. (when the slice is perpendicular to the axis of rotation) But i was going to do the shell method if i could not manage to get the answer using disk/washer

OpenStudy (anonymous):

first is supposed to say disk/washer

OpenStudy (anonymous):

ok here is a picture of my work

OpenStudy (tkhunny):

You can give clues... 1) Which axis are we using as the axis of rotation? 2) Are you using disks/washers or shells? 3) What's in front? \(\pi\;or\;2\pi\) 4) What are the limits? 5) What are your arguments?

OpenStudy (tkhunny):

I ALWAYS do it both ways. This builds great confidence in the result.

OpenStudy (anonymous):

i am rotating about the y-axis using the disk/washer method, pi is in front because it is the area of a circle (the slice we are cutting out of the graph), it is bounded by 0 to 8 on the y-axis, and we use those bc we are using dy

OpenStudy (anonymous):

can you view that?

OpenStudy (tkhunny):

No, sorry.

OpenStudy (anonymous):

ok hold on and i will try to upload it a different way

OpenStudy (tkhunny):

You can just answer my five questions.

OpenStudy (anonymous):

i did. It was above my http link

OpenStudy (anonymous):

try that.

OpenStudy (tkhunny):

Ah! First thing, I read the problem wrong. You have the right region. Forget the answer I gave above. About he y-axis should be \(\dfrac{13312\pi}{9}\)

OpenStudy (anonymous):

it says that the answer is incorrect. I appreciate the effort though. I am just completely lost as to what to do

OpenStudy (tkhunny):

Why on earth would you expand those simple arguments? \(\int (12-y)^{2}\;dy = -\dfrac{(12-y)^{3}}{3} + C\) Use a little Chain Rule, rather than a LOT of algebra. This will prevent errors.

OpenStudy (tkhunny):

Whoops. I mean "27" in the denominator. Not "9". Trouble paying attention, i guess.

OpenStudy (anonymous):

umm because i don't think that you can do integration that way without a u-substitution. I put in 27 and it still says i am incorrect. haha i wasn't joking when i said i had tried just about everything and it denied me

OpenStudy (tkhunny):

\(\int\left(\dfrac{y+4}{3}\right)^{2}\;dy = \dfrac{3}{3}\cdot\left(\dfrac{y+4}{3}\right)^{3} + C\). Call it what you will, Chain Rule or u-substitution, it will make your life easier. Around the y-axis, defintely 13312\(\pi\)/27, assuming we have the irght region. I do believe you picked the right region. You have it right all the way down to the quadratic expression all over 9.

OpenStudy (anonymous):

chain rule is derivatives u sub is for integration

OpenStudy (tkhunny):

Two steps down is no good, though. Your 112 magically turned into 122. p'-tay-to p'-tah-to. Remember when I said not to get stuck in one way of thinking?

OpenStudy (tkhunny):

btw, you appear to have a very good, clear style. Excellent work!

OpenStudy (anonymous):

yeah thanks for catching that. i will look it over. thanks for that compliment and the help!

OpenStudy (tkhunny):

x-axis shoud be \(\dfrac{2048}{9}\cdot\pi\)

OpenStudy (tkhunny):

Shells, y-axis: \(2\pi\int\limits_{4/3}^{4}x\cdot (3x-4)\;dx + 2\pi\int\limits_{4}^{12}x\cdot (12-x)\;dx = \dfrac{13312}{27}\cdot\pi\)

OpenStudy (sirm3d):

 \[\Delta V = \pi(R^2 - r^2)\Delta y\] where \[R=12-y,\quad r=\frac{y+4}{3}\] The volume of the solid of revolution is \[\Large V=\pi\int_{-4}^8\left[ (12-y)^2 - \left(\frac{y+4}{3}\right)^2\right]\,\mathrm dy\]

OpenStudy (sirm3d):